# Microscale architected materials for elastic wave guiding: Fabrication and dynamic characterization across length and time scales

### v1

@@ IV Conclusion
We have presented a new experimental protocol that enables and characterizes elastic wave guiding in free-standing micro-architected materials. Our fabrication protocol adopts state-of-the-art silicon microfabrication techniques to generate free-standing microarchitected films with demonstrated spatial resolutions down to $5$ µm, and unit cell densities on the order of $10^{3}$ per mm ($\sim 6\times 10^{5}$ unit cells on a single free-standing film). These samples support a wide range of elastic wave modes across length scales spanning millimeters down to $10$ µm, and time scales of tens of nanometers to hundreds of microseconds.
@@ Appendix A Pump-probe experiment control
At the start of the experiment, the function generator channels are turned on before running the Python code. Pulses from the AFG are read by the pump laser control unit, which sends read-out signals to the oscilloscope trigger input channel upon triggering each optical pulse. Before each acquisition, the input parameters are sent to the oscilloscope, which is then armed for acquisition. After acquisition, the data is stored locally on the hard drive of the oscilloscope, before triggering one step of the scanning stage. Although the duration for this motion lasts on the order of milliseconds, a wait time of 1 s is imposed to ensure that the scanning stages are stabilized before the next acquisition. This then triggers the next acquisition in a loop, until the maximum number of steps entered a-priori as input is reached. The scanning stage motion was verified to be accurate until $5$ µm, beyond which it was unnecessary to test for the purpose of our experiments. However, this is not the limiting scanning resolution achievable for high-resolution measurements. Future versions of the automation code can involve remote control of the AFG trigger to the pump laser, to limit the number of pump pulses on the sample, and real-time storage and analysis of data on a centralized server. The implementation of these modifications is relatively straightforward, with the exception of requiring large storage space and fast data transfer for high-resolution, long-time measurements. However, such a system could advantageously alter the paradigm of experiment-driven computational design.
@@ Appendix C Finite element analysis
The basis for this analysis is a description of the structure by Timoshenko beam finite elements, modeling a single unit cell. Timoshenko beam elements provide an excellent approximation for the lowest out-of-plane modes as compared to higher-fidelity but numerically inefficient solid finite elements; see [64] for further discussion. Our mesh exhibits a maximum element size of $L/20$, where $L=100\ \mu$m is the lattice spacing, which was found to be sufficiently fine to resolve the first two out-of-plane bands accurately.

### v2

@@ IV Conclusion
We have presented a new experimental protocol that enables and characterizes elastic wave guiding in free-standing micro-architected materials. Our fabrication protocol adopts state-of-the-art silicon microfabrication techniques to generate free-standing microarchitected films with demonstrated spatial resolutions down to 55 µm, and unit cell densities on the order of $10^{3}$ per $m$ ($\sim 6\times 10^{5}$ unit cells on a single free-standing film). These samples support a wide range of elastic wave modes across length scales spanning millimeters down to $10$ µm, and time scales of tens of nanometers to hundreds of microseconds.
@@ Appendix A Pump-probe experiment control
At the start of the experiment, the function generator channels are turned on before running the Python code. Pulses from the AFG are read by the pump laser control unit, which sends read-out signals to the oscilloscope trigger input channel upon triggering each optical pulse. Before each acquisition, the input parameters are sent to the oscilloscope, which is then armed for acquisition. After acquisition, the data is stored locally on the hard drive of the oscilloscope, before triggering one step of the scanning stage. Although the duration for this motion lasts on the order of milliseconds, a wait time of 1 s is imposed to ensure that the scanning stages are stabilized before the next acquisition. This then triggers the next acquisition in a loop, until the maximum number of steps entered a-priori as input is reached. The scanning stage motion was verified to be accurate until 55 µm, beyond which it was unnecessary to test for the purpose of our experiments. However, this is not the limiting scanning resolution achievable for high-resolution measurements. Future versions of the automation code can involve remote control of the AFG trigger to the pump laser, to limit the number of pump pulses on the sample, and real-time storage and analysis of data on a centralized server. The implementation of these modifications is relatively straightforward, with the exception of requiring large storage space and fast data transfer for high-resolution, long-time measurements. However, such a system could advantageously alter the paradigm of experiment-driven computational design.
@@ Appendix C Finite element analysis
[…]